OpenStudy (anonymous):

Two integers, A and B, have a product of -48. What is the greatest possible sum of A and B?

jimthompson5910 (jim_thompson5910):

"Two integers, A and B, have a product of -48" which means A*B = -48

jimthompson5910 (jim_thompson5910):

I'm going to solve for B to get B = -48/A I just divided both sides by A

jimthompson5910 (jim_thompson5910):

now we're asked this: What is the greatest possible sum of A and B? Let y = A+B then plug in B = -48/A to get y = A + (-48/A) and finally replace A with x. So we're making A = x y = x + (-48/x) with me so far?

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

well this isn't finished though, but hopefully you see how to finish up

OpenStudy (anonymous):

Basically one of ur numbers will be negative. 6*-8= -48